OpenStudy (anonymous):

i need to find the positive and negative coterminal angels for 450 degrees and 11pi/6

OpenStudy (anonymous):

11pi/6 + 2pi = 11pi/6 + 12pi/6 = 23pi/6

OpenStudy (anonymous):

450=360+90=90 degree. Again 450=720-270=-270 degree \[\frac{ 11\pi }{6 }=2\pi-\frac{ \pi }{ 6 }=-\frac{ \pi }{6 }~radians.\]